# We Have a Package for You! A Comprehensive Analysis of Package Hallucinations by Code Generating LLMs

### v1

@@ 6.2 Mitigation Implementation and Results
Retrieval Augmented Generation (RAG). We utilized a method of before generation RAG to supplement the prompt with valid package names to assist the model in generating a response. We made an additional dataset to serve as the supplementary information by taking the top 20,000 most popular packages from PyPI and asking LlaMA-2 to generate a list of 5 questions that each package could help answer. After removing duplicate responses, this resulted in 65,000 statements in the form “Package [x] could answer questions about [y]”. These 65,000 statements were stored in a vector database, which enables efficient retrieval of semantically similar statements. When a model was asked to recommend packages given a code generation prompt or Stack Overflow question, the vector database would first return the top 5 most semantically similar statements. These statements are prepended to the prompt to give the model additional information containing established valid packages to generate a non-hallucinated response.
@@
Self-Refinement. Drawing on insights from our findings in RQ3 (see Section 5.3), which revealed that LLMs often exhibit proficiency in identifying their own package hallucinations, we implemented a self-refinement method. Following the generation of package names, the model is queried regarding the validity of these packages. If the model indicates that the packages are invalid, the response is regenerated with a specific instruction to not use the invalid package. This regeneration process is allowed to iterate up to five times, acknowledging that many package hallucinations are persistent, as demonstrated in RQ3, and may be repeatedly generated.
@@
Fine-tuning. For our next method, we fine-tuned the models using the code/package list (Heuristic 1) and prompt/package list (Heuristic 2) pairs that were generated during our initial experiments. (Section 5.1) All hallucinations were filtered out and the models were re-trained using the remaining valid responses (560,000 samples).
@@
We implemented these mitigation techniques using the DeepSeek Coder 6B and CodeLlama 7B models. These models were selected because they represent two distinct classes of foundational models, with these specific parameter sizes reflecting diverse performance levels: DeepSeek being among the best-performing, and CodeLlama among the worst-performing during our initial experiments (see Section 5.1). Both models were tested using each of the above methods individually and then using all three methods in an ensemble configuration.
TABLE II: Performance of the mitigation techniques.
@@
DeepSeek CodeLlama Baseline (No Mitigations) 16.14% 26.28% Retrieval Augmented Generation (RAG) 12.24% 13.40% Self-Detected Feedback 13.04% 25.51% Fine-tuning 2.66% 10.27% Ensemble 2.40% 9.32%
@@
Results. Overall, all the mitigation strategies we implemented resulted in a reduced rate of package hallucination, with RAG and Supervised Fine-Tuning proving particularly effective (see Table II). Fine-tuning proved to significantly improve results, especially for the DeepSeek model, where hallucinations were reduced by 83%, achieving a total rate of just 2.66%, which is a lower rate than any of the ChatGPT models (observed during RQ1). Self-detected feedback was also much more effective for the DeepSeek model (19% reduction) compared to the CodeLlama model (3% reduction). This aligns with our results in RQ3, where the DeepSeek model was proficient at detecting hallucinations while the CodeLlama model had a strong bias towards labeling packages as valid, that limited its ability to reliably detect errors. The ensemble method of combining all mitigation strategies further improved results, reducing hallucination rates by 85% and 64% from their baseline levels for DeepSeek and CodeLlama, respectively.
@@ Appendix A Truncated List of LLM Generated Coding Prompts
  3. 3.
    Generate Python code that implements a simple flock-based file locking mechanism using the ‘with‘ syntax, allowing multiple threads to safely access a shared resource while ensuring exclusive access for writing operations.
  4. 4.
    Generate Python code that implements a backport of f-strings to Python versions prior to 3.6 using the ‘str.format()‘ method.
  5. 5.
@@
  6. 6.
    Generate Python code that implements a rate limiter for Flask applications using the ‘limiter‘ library, which provides a simple way to add rate limiting to any Flask endpoint.
@@ Appendix B System Messages and Prompts
Figures 8, 9 and 10 below show the system messages and prompts that are sent to each model to generate the code samples and package lists.
@@
[LLM Generated prompt from Appendix A or Stack Overflow question.
@@
Which [Python/JavaScript] packages are required to run this code: [Code sample generated from code generation phase]
@@
Which [Python/JavaScript] packages would be useful in solving the following coding problem: [Original LLM generated prompt or Stack Overflow question]
@@
Appendix C Model Parameters
Table III shows the model parameters used during our RQ1 tests.
TABLE III: An overview of the model parameters.
@@ Appendix C Model Parameters
Appendix E Cross Language Hallucinations
Table IV displays the number of packages that were generated for Python specific code and were detected as hallucinations, but are actually valid packages in other programming languages.
TABLE IV: Confusion by programming language repository.
Appendix F Hallucination Detection Performance
Table V shows average precision and recall values for our test on LLMs’ hallucination detection performance (as described in Section 5.3).
TABLE V: Performance of hallucination detection tests.
Appendix G Complete Results for Python and JavaScript
Table VI and Table VII shows complete results for package hallucination experiments observed across all tested models for Python and JavaScript.
TABLE VI: Hallucination Percentages for all models tested using Python code
Model Total Hallucination LLM Generated Prompts Stack Overflow Prompts “pip install” GPT-4 Turbo 3.59% (2,739/76,313) 3.29% (1,518/46,204) 4.07% (1,169/28,728) 3.77% (52/1,381) GPT-4 4.05% (2,969/73,396) 3.83% (1,741/45,403) 4.45% (1,046/23,487) 4.04% 182/4,506 GPT-3.5 Turbo 5.76% (4,387/76,123) 5.98% (2,495/41,7) 5.50% (1,868/33,948) 5.63% (24/426) DeepSeek 1B 13.63% (12,481/91,543) 11.07% (5,847/52,806) 16.39% (5,901/36,007) 26.85% (733/2,730) DeepSeek 33B 16.53% (7,071/42,788) 13.85% (3,623/26,167) 25.47% (3,033/11,906) 8.80% (415/4,715) WizardCoder 33B 14.31% (4,909/34,300) 9.79% (1,579/16,125) 21.40% (2,8523/13,329) 9.84% (477/4,846) DeepSeek 6B 16.61% (16,526/99,505) 14.01% (9,240/65,957) 23.56% (6,792/28,828) 10.47% (494/4,720) OpenChat 7B 18.31% (16,932/92,452) 17.39% (9,582/55,092) 19.98% (6,454/32,307) 17.73% (896/5,053) CodeLlama 13B 18.03% (12,404/68,809) 15.21% (6,450/42,410) 22.76% (5,752/25,273) 17.94% (202/1,126) Mixtral 8x7B 16.79% (7,753/46,166) 13.12% (2,749/20,951) 20.92% (4,068/19,949) 16.23% (936/5,766) MagiCoder 7B 16.60% (20,258/122,057) 15.76% (11,994/76,096) 18.48% (7,621/41,248) 13.64% (643/4,713) CodeLlama 34B 21.15% (24,905/117,777) 15.22% (9,495/62,366) 28.56% (14,891/52,135) 15.84% (519/3,276) Mistral 7B 20.71% (7,959/38,437) 14.47% (2,808/19,412) 30.69% (3,922/12,778) 19.67% (1,229/6,247) WizardCoder 7B - Python 20.69% (11,408/55,131) 16.80% (4,698/27,962) 26.73% (6,112/22,867) 13.90% (598/4,302) CodeLlama 34B - Python 20.97% (12,128/57,833) 19.01% (5,913/31,112) 23.39% (6,208/26,540) 3.87% (7/181) CodeLlama 7B 26.12% (27,814/106,487) 21.51% (12,961/60,261) 32.53% (14,671/45,099) 16.15% (182/1,127)
TABLE VII: Hallucination percentages for all models tested using JavaScript code

### v2

@@ 6.2 Mitigation Implementation and Results
Retrieval Augmented Generation (RAG). We utilized a method of before generation RAG to supplement the prompt with valid package names to assist the model in generating a response. We made an additional dataset to serve as the supplementary information by taking the top 20,000 most popular packages from PyPI and prompting LlaMA-2 to generate a list of 5 questions that each package could help answer given the description. After removing duplicate responses, this resulted in 65,000 statements in the form “Package [x] could answer questions about [y]”. These 65,000 statements were stored in a vector database, which enables efficient retrieval of semantically similar statements. When a model was asked to recommend packages given a code generation prompt or Stack Overflow question, the vector database is first queried from within the code to return the top 5 most semantically similar statements. These statements are appended to the prompt to give the model additional information containing established valid packages to generate a non-hallucinated response.
@@
Self-Refinement. Drawing on insights from our findings in RQ3 (see Section 5.3), which revealed that LLMs often exhibit proficiency in identifying their own package hallucinations, we implemented a self-refinement method. Following the generation of package names, the model is queried regarding the validity of these packages. If the model indicates that the packages are invalid, the response is regenerated with a specific instruction to not use the invalid package. This regeneration process is allowed to iterate up to five times, acknowledging that many package hallucinations are persistent, as demonstrated in RQ3, and may be repeatedly generated. It is possible during this test that a valid package is misclassified as a hallucination by the model, although because this is an iterative process the success rate should outweigh the false positive rate over the course of testing, resulting in a net decrease in hallucination rate.
@@
Fine-tuning. For our next method, we fine-tuned the models using the code/package list (Heuristic 1) and prompt/package list (Heuristic 2) pairs that were generated during our initial experiments (Section 5.1). All hallucinations were filtered out and the models were re-trained using the remaining valid responses (560,000 samples). As fine-tuning affects the underlying model weights, we also need to make sure the fine-tuned model retains the ability to produce functional and effective code, which we will test by comparing the code quality from the original models and the fine-tuned models using the well-known HumanEval benchmark [8]. The HumalEval benchmark is a set of pre-defined prompts and test cases and the final score is a percentage of problems for which the model generated code is both syntactically correct and passes the test cases, reflecting the model’s ability to produce functional and accurate code.
@@
We implemented these mitigation techniques using the DeepSeek Coder 6.7B and CodeLlama 7B models. These models were selected because they represent two distinct classes of foundational models, with these specific parameter sizes reflecting diverse performance levels: DeepSeek being among the best-performing, and CodeLlama among the worst-performing during our initial experiments (see Section 5.1). Both models were tested using each of the above methods individually and then using all three methods in an ensemble configuration.
Table 4: Performance of the mitigation techniques.
@@
DeepSeek CodeLlama Baseline (No Mitigations) 16.14% 26.28% Retrieval Augmented Generation (RAG) 12.24% 13.40% Self-Refinement 13.04% 25.51% Fine-tuning 2.66% 10.27% Ensemble 2.40% 9.32%
@@
Results. Overall, all the mitigation strategies we implemented resulted in a reduced rate of package hallucination, with RAG and Supervised Fine-Tuning proving particularly effective (see Table 4). Fine-tuning proved to significantly improve results, especially for the DeepSeek model, where hallucinations were reduced by 83%, achieving a total rate of just 2.66%, which is a lower rate than any of the ChatGPT models (observed during RQ1). Self-refinement feedback was also much more effective for the DeepSeek model (19% reduction) compared to the CodeLlama model (3% reduction). This aligns with our results in RQ3, where the DeepSeek model was proficient at detecting hallucinations while the CodeLlama model had a strong bias towards labeling packages as valid, that limited its ability to reliably detect errors. The ensemble method of combining all mitigation strategies further improved results, reducing hallucination rates by 85% and 64% from their baseline levels for DeepSeek and CodeLlama, respectively.
The code quality of the fine-tuned models did decrease significantly, -26.1% and -3.1% for DeepSeek and CodeLlama respectively, in exchange for substantial improvements in package hallucination rate. While the code quality was negatively impacted, the fine-tuned scores are still at comparable levels to other high performing models such as Mistral 7B (26.1%), Llama 65B (23.1%), and Llama 2 7B (12.8%).
Table 5: Analysis of code quality after fine-tuning.
DeepSeek CodeLlama Original Model pass@1 51.4% 19.6% Fine-tuned pass@1 25.3% 16.4%
@@ 7 Discussion and Conclusion
Ethics Considerations
We have disclosed our research to model providers including OpenAI, Meta, DeepSeek, and Mistral AI. As of this writing we have received no response or feedback. Our research highlights the feasibility of a new attack vector that can be used to carry out package confusion attacks by exploiting the package hallucinations generated by code-generating LLMs. Hallucinations are a well-studied limitation of generative AI models, including code-generating LLMs, which even current state-of-the-art techniques cannot completely eliminate. Our goal in this research is to get a better understanding of this phenomenon of package hallucinations in code-generating LLMs. We hope that these findings can help secure future models so that users are better protected against package confusion threats enabled due to package hallucinations.
One course of action that we chose not to pursue for ethical reasons was publishing actual packages using hallucinated package names to PyPI (with no actual code). There would be scientific value in demonstrating the validity of exploiting package hallucinations on an open-source repository, but we felt that doing so would have been misleading, undermined the integrity of the repository, wasted resources, and could be interpreted as violating the terms of service.
No human subjects were involved in our research and all experiments were conducted in controlled settings with no impact on external persons or entities.
We are committed to the open science policy and will make all source code, datasets, and generated code publicly available. The only exception is we will not release our master list of hallucinated package names to the public. Such information could be misused by malicious actors to execute a package hallucination attack. The master list will be shared responsibly by request to verified researchers.
Another consideration for making our data public is that half of our prompt dataset was compiled using user created data from StackOverflow. This data consisted of carefully scraped questions and the only data obtained was the question itself; we did not collect any user information or even the answers to the question. This approach aligns with privacy best practices and adheres to data minimization principles, which dictate collecting only the data necessary for our research objectives. Furthermore, the process of scraping was designed to comply with StackOverflow’s terms of service and data use policies, ensuring that we maintained legal and ethical integrity throughout the data collection phase. This careful consideration in data handling not only safeguards user privacy but also reinforces the ethical standards we uphold in making our dataset publicly available.
@@ Appendix A Truncated List of LLM Generated Coding Prompts
  3. 3)
    Generate Python code that implements a simple flock-based file locking mechanism using the ‘with’ syntax, allowing multiple threads to safely access a shared resource while ensuring exclusive access for writing operations.
  4. 4)
    Generate Python code that implements a backport of f-strings to Python versions prior to 3.6 using the ‘str.format()’ method.
  5. 5)
@@
  6. 6)
    Generate Python code that implements a rate limiter for Flask applications using the ‘limiter’ library, which provides a simple way to add rate limiting to any Flask endpoint.
@@ Appendix B System Messages and Prompts
Figures 10, 11 and 12 below show the system messages and prompts that are sent to each model to generate the code samples and package lists.
@@
LLM Generated prompt from Appendix A or Stack Overflow question.
@@
Python/JavaScript] packages are required to run this code: [Code sample generated from code generation phase].
@@
Which [Python/JavaScript] packages would be useful in solving the following coding problem: [Original LLM generated prompt or Stack Overflow question].
@@
System Message
You are a coding assistant that assists users in creating simple prompts that will be used to generate [Python/JavaScript] code. No code should be used in the response.
Prompt
Your answer must begin with ‘Generate [Python/JavaScript] code that’ and must not be longer than one sentence. Do not include extra text or formatting (i.e. do not start with ‘Sure! Here’s a prompt…’). Write a prompt that would generate [Python/JavaScript] code to accomplish the same tasks as the following package description: [package description from PyPI].
Figure 13: Coding prompt generation.
Appendix C Model Parameters and Testing Environment
Model testing was conducted in two distinct computing environments - a Debian environment with 40 nodes, each equipped with 40 CPU cores, 1TB of RAM, and NVIDIA A100 or V100 GPUs and a Ubuntu system with 80 CPU cores, 750 GB of RAM, and 3 NVIDIA RTX 6000 GPUs.
Table 6 shows the model parameters used during our RQ1 tests.
Table 6: An overview of the model parameters.
@@ Appendix C Model Parameters and Testing Environment
Appendix E Complete Results for Python and JavaScript
Table 7 and Table 8 shows complete results for package hallucination experiments observed across all tested models for Python and JavaScript.
Table 7: Hallucination Percentages for all models tested using Python code.
Model Total Hallucination LLM Generated Prompts Stack Overflow Prompts “pip install" GPT-4 Turbo 3.59% (2,739/76,313) 3.29% (1,518/46,204) 4.07% (1,169/28,728) 3.77% (52/1,381) GPT-4 4.05% (2,969/73,396) 3.83% (1,741/45,403) 4.45% (1,046/23,487) 4.04% 182/4,506 GPT-3.5 Turbo 5.76% (4,387/76,123) 5.98% (2,495/41,7) 5.50% (1,868/33,948) 5.63% (24/426) DeepSeek 1B 13.63% (12,481/91,543) 11.07% (5,847/52,806) 16.39% (5,901/36,007) 26.85% (733/2,730) DeepSeek 33B 16.53% (7,071/42,788) 13.85% (3,623/26,167) 25.47% (3,033/11,906) 8.80% (415/4,715) WizardCoder 33B 14.31% (4,909/34,300) 9.79% (1,579/16,125) 21.40% (2,8523/13,329) 9.84% (477/4,846) DeepSeek 6B 16.61% (16,526/99,505) 14.01% (9,240/65,957) 23.56% (6,792/28,828) 10.47% (494/4,720) OpenChat 7B 18.31% (16,932/92,452) 17.39% (9,582/55,092) 19.98% (6,454/32,307) 17.73% (896/5,053) CodeLlama 13B 18.03% (12,404/68,809) 15.21% (6,450/42,410) 22.76% (5,752/25,273) 17.94% (202/1,126) Mixtral 8x7B 16.79% (7,753/46,166) 13.12% (2,749/20,951) 20.92% (4,068/19,949) 16.23% (936/5,766) MagiCoder 7B 16.60% (20,258/122,057) 15.76% (11,994/76,096) 18.48% (7,621/41,248) 13.64% (643/4,713) CodeLlama 34B 21.15% (24,905/117,777) 15.22% (9,495/62,366) 28.56% (14,891/52,135) 15.84% (519/3,276) Mistral 7B 20.71% (7,959/38,437) 14.47% (2,808/19,412) 30.69% (3,922/12,778) 19.67% (1,229/6,247) WizardCoder 7B - Python 20.69% (11,408/55,131) 16.80% (4,698/27,962) 26.73% (6,112/22,867) 13.90% (598/4,302) CodeLlama 34B - Python 20.97% (12,128/57,833) 19.01% (5,913/31,112) 23.39% (6,208/26,540) 3.87% (7/181) CodeLlama 7B 26.12% (27,814/106,487) 21.51% (12,961/60,261) 32.53% (14,671/45,099) 16.15% (182/1,127)
Table 8: Hallucination percentages for all models tested using JavaScript code.
@@ Appendix E Complete Results for Python and JavaScript
Appendix F Additional Background Information
F.1 Large Language Models
The advent of extremely versatile and high-performing attention-based transformer models [75] have catalyzed the emergence of foundational language models, which are large-scale models described by billions of parameters and pre-trained on vast datasets. These foundational LLMs can be tailored to perform a wide range of downstream tasks without the need for task-specific architecture modifications [4]. Models such as BERT (Bidirectional Encoder Representations from Transformers) [10], CLIP (Contrastive Language-Image Pretraining) [63], and GPT (Generative Pre-Trained Transformer) [5] are examples of foundational language models that have emerged as new benchmarks for a wide range of tasks.
F.2 Hallucinations
The persistent issue of hallucinations in LLMs has spurred extensive research into various mitigation strategies for standard natural language generation tasks, broadly categorized into prompt engineering and model architecture enhancements [73]. Prompt engineering techniques such as Retrieval Augmented Generation (RAG) [37] and self-refinement methods [50] aim to refine the input provided to the model to produce more accurate outputs. Alternatively, developing more robust models involves approaches such as supervised fine-tuning [71], inference time intervention [40], and incorporating knowledge graphs [25] to improve model understanding and reduce errors/hallucinations in the model output.
The manipulation of the temperature parameter within each model has also been shown to significantly influence the prevalence of hallucinations in LLM output [2]. The temperature parameter modulates the probability distribution over potential output tokens [53]. Lower temperatures produce less random, more predictable outputs, while higher temperatures increase the likelihood of sampling low-frequency tokens, raising the risk of hallucinations [11].
Appendix G Packages and Modules Explanation
Here is a quick primer on the purpose of packages and modules in coding languages and how that relates to the topic of package hallucinations. To enable code modularity and reusability, interpreted programming languages such as Python and JavaScript allow for entities called modules. A module is a chunk of code, often in an external file, that performs a specific task or function. By encapsulating related code into modules, developers can organize their programs more efficiently and make them easier to maintain and share. A package is a collection of related modules that work together to provide certain functionality. To use a particular module in their source code, a developer must install the appropriate package into its development environment by first downloading the package from an online package manager or repository such as PyPI or npm, if it is not locally available, and then import the desired module into the code using appropriate import functions, e.g., import in Python or require in JavaScript.
These module names do not necessarily need to match the package names and the namespace for modules is not protected, i.e., different packages may include modules of the same name. This discrepancy poses a significant challenge for detecting package dependencies from the raw Python and JavaScript code, as the import/require statements that are typically included in code samples for importing modules do not have a unique mapping to package names. Thus, it would not be accurate to simply parse the generated code for import or require statements, as those statements refer to module names and not package names.
Appendix H Stack Overflow Tags
These are the "tags" used to filter questions on StackOverflow to ensure that all questions contained relevant subject matter for the respective programming language. Some tags are relevant to all programming and thus appear in both sets of tags, while others are language specific. We considered any tag with more than 5,000 questions and manually selected the applicable topics, resulting in 233 tags for Python and 274 for JavaScript
H.1 Python Tags
python, arrays, django, pandas, string, dataframe, multithreading, list, algorithm, image, forms, numpy, function, loops, rest, apache, csv, dictionary, unit-testing, tensorflow, file, class, sorting, date, authentication, opencv, matplotlib, http, validation, sockets, object, xaml, oop, if-statement, web-services, email, kubernetes, github, parsing, user-interface, pointers, security, machine-learning, flask, exception, debugging, tkinter, ssl, web-scraping, testing, recursion, animation, session, math, django-models, keras, image-processing, dom, logging, matrix, post, pyspark, optimization, networking, opengl, events, encryption, audio, canvas, indexing, multidimensional-array, random, vector, video, data-structures, serialization, model-view-controller, iframe, beautifulsoup, design-patterns, text, mobile, lambda, graph-theory, websocket, scikit-learn, file-upload, deep-learning, error-handling, charts, server, https, merge, jupyter-notebook, encoding, django-views, hash, automation, pytorch, utf-8, python-requests, scipy, file-io, neural-network, interface, linked-list, django-templates, functional-programming, pygame, scrapy, io, bluetooth, 3d, dns, pyqt, package, tree, statistics, computer-vision, dockerfile, hashmap, character-encoding, cryptography, pyqt5, iterator, multiprocessing, cuda, ftp, plotly, operating-system, kivy, type-conversion, dependencies, geometry, format, subprocess, swagger, mapreduce, widget, set, tuples, smtp, ldap, return, dataset, network-programming, base64, stack, iteration, embedded, udp, gmail, conv-neural-network, bots, django-admin, time-complexity, pytest, python-imaging-library, web-crawler, hex, substring, filesystems, export, jinja2, seaborn, regression, thread-safety, logic, cloud, compression, data-science, celery, gpu, pandas-groupby, artificial-intelligence, video-streaming, backend, client, conda, jupyter, refactoring, bigdata, bit-manipulation, classification, boto3, schema, 2d, data-visualization, python-asyncio, wxpython, ipython, nltk, shader, rendering, pymongo, django-queryset, binary-tree, storage, blockchain, ethereum, vectorization, solidity, heap-memory, pyinstaller, linear-regression, metadata, kibana, cluster-analysis, networkx, driver, html-parsing, python-import, sympy, dynamic-programming, listener, pyqt4, user-input, cluster-computing, openpyxl, lxml, cython, simulation, spyder, decorator, pickle, python-multiprocessing, python-sphinx, python-multithreading, ironpython, python-tesseract, python-telegram-bot, python-polars, python-xarray
H.2 JavaScript Tags
javascript, arrays, jquery, reactjs, node.js, angularjs, typescript, string, database, react-native, string, multithreading, list, algorithm, image, function, vue.js, loops, rest, apache, csv, dictionary, unit-testing, file, class, sorting, date, authentication, opencv, http, validation, sockets, object, xaml, oop, if-statement, web-services, email, kubernetes, github, parsing, user-interface, selenium-webdriver, pointers, security, machine-learning, listview, exception, debugging, ssl, web-scraping, testing, npm, mongoose, recursion, animation, session, math, webpack, heroku, redirect, dom, image-processing, logging, matrix, post, button, next.js, jquery-ui, optimization, d3.js, caching, networking, opengl, events, encryption, audio, canvas, indexing, multidimensional-array, redux, cookies, random, vector, video, data-structures, serialization, model-view-controller, iframe, design-patterns, text, mobile, react-hooks, google-chrome-extension, meteor, bootstrap-4, graph-theory, websocket, file-upload, deep-learning, error-handling, browser, vuejs2, charts, view, highcharts, server, https, merge, google-maps-api-3, extjs, axios, encoding, ember.js, hash, automation, react-redux, utf-8, jestjs, neo4j, jquery-mobile, material-ui, file-io, interface, linked-list, functional-programming, three.js, rxjs, datatables, backbone.js, react-router, knockout.js, angular-material, jwt, clojure, angularjs-directive, io, discord.js, kendo-ui, bluetooth, 3d, dns, package, tree, electron, frontend, dockerfile, hashmap, character-encoding, cryptography, iterator, multiprocessing, ftp, plotly, operating-system, jasmine, vuejs3, gulp, nuxt.js, dom-events, sequelize.js, nestjs, type-conversion, dependencies, geometry, webdriver, chart.js, format, swagger, widget, set, xmlhttprequest, tuples, smtp, ldap, return, dataset, network-programming, base64, protractor, arguments, stack, iteration, embedded, udp, gmail, bots, time-complexity, fetch-api, local-storage, angular-ui-router, cypress, coffeescript, babeljs, web-crawler, hex, log4j, substring, filesystems, export, regression, thread-safety, logic, cloud, dojo, jsx, gruntjs, vuetify.js, angularjs-ng-repeat, angularjs-scope, compression, video-streaming, backend, client, refactoring, bigdata, bit-manipulation, classification, mocha.js, momentjs, vuex, requirejs, bootstrap-modal, handlebars.js, phantomjs, angular-cli, passport.js, javascript-objects, eslint, lodash, phpstorm, puppeteer, schema, 2d, data-visualization, shader, rendering, binary-tree, storage, blockchain, visualization, ethereum, vectorization, ejs, sails.js, extjs4, vue-router, this, pug, node-modules, solidity, heap-memory, metadata, svelte, kibana, driver, underscore.js, office-js, ember-data, webstorm, npm-install, html-parsing, dynamic-programming, settimeout, facebook-javascript-sdk, listener, user-input, cluster-computing, lxml, openlayers, simulation, pickle, google-tag-manager, javascript-objects